Four Finger Appliance Company manufactures small kitchen appliances. The product line consists of blenders and toaster ovens. Four Finger Appliance presently uses the multiple production department factory overhead rate method. The factory overhead is as follows:

1

Assembly Department

$275,736.00

2

Test and Pack Department

201,046.00

3

Total

$476,782.00

The direct labor information for the production of 8,390 units of each product is as follows:

Department
Assembly Test and Pack
Blender 740 dlh 2,040 dlh
Toaster oven 2,205 680
Total 2,945 dlh 2,720 dlh

Four Finger Appliance used direct labor hours to allocate production department factory overhead to products.

The management of Four Finger Appliance Company has asked you to use activity-based costing instead of direct labor hours to allocate factory overhead costs to the two products. You have determined that $81,366 of factory overhead from each of the production departments can be associated with setup activity ($162,732 in total). Company records indicate that blenders required 142 setups, while the toaster ovens required only 49 setups. Each product has a production volume of 8,390 units.

Required:
Complete the Activity Tables for blenders and toaster ovens.
a. Determine the three activity rates (assembly, test and pack, and setup).*
b. Determine the total factory overhead and factory overhead per unit allocated to each product using the activity rates in (a).*
*If required, round all per-unit answers to the nearest cent.
Answer #1

a. The three activity rates (assembly, test and pack, and setup)

Cost pools Cost Cost drivers Activity Rates per unit
Assembly 194370 2945 dlh                                    66
Test and pack 119680 2720 dlh                                    44
Setup 162732 191 Setups (142+49)                                 852

b.

Calculation of over head rates per unit as per ABC
Blender Toaster oven
(i) (ii)= (i)*cost driver (iii)= (ii)/no. of units (ii)= (i)*cost driver (iii)= (ii)/no. of units
Activity Activity rate per unit total overhead cost Over head Rate per unit total overhead cost Over head Rate per unit
Assembly 66 48840 (66*740) 5.82 145530 (66*2205) 17.35
Test and pack 44 89760 (44*2040) 10.70 29920 (44*680) 3.57
Setup 852 120984 (852*142) 14.42 41748 (852*49) 4.98